OpenStudy (anonymous):

An original rectangle has a length of fourteen and a width of twelve. What happens to the area if the new width is three times the original width?

OpenStudy (anonymous):

Recall the area of the rectangle. A = lw A = (14)(12) A = 168 units² Tripled... A = l(3w) A = (14)(3)(12) A = 504 units³ If you triple the width, then you also triple the area because they are directly proportional to each other.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No Problem :)